# MoodLoopGP: Generating Emotion-Conditioned Loop Tablature Music with Multi-Granular Features

### 23 Jan 2024 v1

@@ 3.3 Bar-level Features
We utilize tonal tension (i.e., cloud diameter, cloud momentum, and tensile strain) as the bar-level features, employing the midi-miner package [26] for feature calculation from musical scores. Similarly, we discretized those values to use them as control tokens. We discretized bar-level features into four levels, using the first quartile, median, and third quartile of the data distribution as separating thresholds. The reason we use four levels instead of two levels to represent bar-level features is that they can create more combinations. Since the bar-level features are added for each bar of the music, they represent the “state” of that bar, and their purpose is to guide the music generation process. Hence, we would want the possible combinations of the features to be relatively large, so that they can represent more kinds of music. In this study, we used three bar-level features. If these features had two levels, there would be a total of 8 possible combinations. However, if the features had four levels, the number of combinations increases to 64.
@@ 5.1 Emotion Identification
The results in Table 2 demonstrate that all the features are important for achieving the best performance. When any of the features are removed, the performance drops significantly. It should be highlighted that when the Emotion labels are missing, the HAP and MAS drop by roughly 30%, and the HVP and MVS drop the most when the Music Psychology Features are missing. This illustrates that the Emotion Labels contribute more to the arousal and Music Psychology Features contribute more to the valence. Additionally, tonal tension seems to contribute more to the valence than arousal, which is evident by the fact that both the HAP and MAS score between All and Missing TT group are roughly the same.
@@ 5.2 Loop Extraction
Table 3 shows the loop extraction evaluation result. MoodLoopGP can generate 45% more loops than the baseline, which demonstrates the advantage of the loop extraction algorithm is successfully kept in MoodLoopGP. We also performed a Wilcoxon Signed-Rank Test to examine the difference between MoodLoopGP and the baseline model. The result (statistic = 2990.0, p < 1e-40) shows that there is a significant effect of the model type on the number of loops generated.
@@ 5.3 Subjective Evaluation
Table 6: Results for all the Likert scale questions, including the listener’s preference, loop coherence (LC), Positive Emotion Scores (PES), and Negative Emotion Scores (NES).
Average Score Preference LC PES NES
@@
Figure 1: Boxplot of the Likert-Scale Questions among different groups.
Table 6 shows the mean scores for Questions 3 to 5 on a 7-point Likert scale. The left-most answer is assigned -3, the right-most answer is assigned 3, and the stride is 1. This is to place the neutral answer (i.e., 0.) in the middle so that positive mean scores indicate positive ratings from the participants. Figure 1 shows the boxplot of the Likert-scale questions. Human-composed music consistently outperforms machine-composed music, indicating the gaps between human-composed and machine-composed music. Loop coherence scores are positive but close to 0 for all generated music groups, indicating loop coherence to listeners is not very strong. However, human and machine compositions have similar loop coherence scores, indicating the loops by the machine are decent. For the emotions conveyed by the music, the human group achieves the best result. The machine-composed group also shows that listeners can recognize the emotional difference between happy and sad music. The boxplot highlights that participants had difficulty differentiating happy and sad music in the "Without Tonal Tension" group, but were able to do so in the "All Information" group. Therefore, Tonal Tension likely helped in generating human-perceivable happy and sad music. This is supported by the Wilcoxon Signed Rank Test results, which will be covered later.
Table 7: Kruskal-Wallis H-test result for the three groups of music in the listening test.
Question H(2) p value
Preference 51.36 7.05e-12
Loop 30.02 3.02e-7
Emotion 21.68 1.96e-5
@@
In order to gain a better understanding of the outcomes of the Likert-scale questions, we conducted a Kruskal-Wallis H-test among the three groups of music. The results, as presented in Table 7, indicate that the source of generation (human or machine) has a significant impact on the listener’s preference, loop, and emotional perspective, as all three questions have a p-value much smaller than $\alpha$ (p < 1e-4). We also carried out a Wilcoxon Sign Rank Test between the Human group and the All Information group for pairwise comparison and found that the small p-values (p < 0.01) demonstrate the significance in all the questions between the groups. Additionally, we were interested in exploring the effect of bar-level features (i.e. Tonal Tension) in the generation process, so we conducted another Wilcoxon Sign Rank Test between the All Information group and the Without Tonal Tension group. Surprisingly, we found that only the p-value of the emotion question is small enough to indicate that Tonal Tension did not enhance the model’s ability to produce high-quality music or coherent loops. However, as previously discovered, it did contribute to generating emotion-specific music.

### 25 Jan 2024 v2

@@ 3.3 Bar-level Features
We utilize tonal tension (i.e., cloud diameter, cloud momentum, and tensile strain) as the bar-level features, employing the midi-miner package [26] for feature calculation from musical scores. Similarly, we discretized those values to use them as control tokens. We discretized bar-level features into four levels, using the first quartile, median, and third quartile of the data distribution as separating thresholds. The reason we use four levels instead of two levels to represent bar-level features is to support more combinations. Since the bar-level features are added for each bar of the music, they represent the “state” of that bar, and their purpose is to guide the music generation process. Hence, we would want the number of possible combinations of the features to be relatively large, so that they can represent more creative possibilities. In this study, we used three bar-level features. If these features had two levels, there would be a total of 8 possible combinations. However, if the features had four levels, the number of combinations increases to 64.
@@ 5.1 Emotion Identification
The results in Table 2 demonstrate that all the features are important for achieving the best performance. When any of the features are removed, the performance drops significantly. It should be highlighted that when the Emotion labels are missing, the HAP and MAS drop by roughly 30%, and the HVP and MVS drop the most when the Music Psychology Features are missing. This illustrates that the Emotion Labels seem to contribute more to the arousal and Music Psychology Features seem to contribute more to the valence. Additionally, tonal tension seems to contribute more to the valence than arousal, as both the HAP and MAS Happy/Sad scores between the All and Missing TT groups are roughly the same, whereas relatively large differences are obtained for the HVP and MVS Happy/Sad scores. Removing tonal tension yields the highest difference between Happy and Sad for MAS, however it is close to the difference obtained when all features are used.
@@ 5.2 Loop Extraction
Table 3 shows the loop extraction evaluation result. MoodLoopGP can generate 45% more loops than the baseline, which demonstrates the advantage of the loop extraction algorithm is successfully kept in MoodLoopGP. We also performed a Wilcoxon Signed-Rank Test to examine the difference between MoodLoopGP and the baseline model. The result (Z = 2990.0, p < 1e-40) shows that there is a significant effect of the model type on the number of loops generated.
@@ 5.3 Subjective Evaluation
Table 6: Results for all the Likert scale questions, including the listener’s preference, loop coherence (LC), Happy Emotion Scores (HES), and Sad Emotion Scores (SES).
Average Score Preference LC HES SES
@@
Figure 1: Boxplot of the Likert-Scale Questions among different groups. The Happy Emotion and Sad Emotion groups correspond to pieces classified as happy and sad, respectively.
Table 6 shows the mean scores for Questions 3 to 5 on a 7-point Likert scale. The left-most answer is assigned -3, the right-most answer is assigned 3, and the stride is 1. This is to place the neutral answer (i.e., 0.) in the middle so that positive mean scores indicate positive ratings from the participants. Figure 1 shows the boxplot of the Likert-scale questions. Human-composed music consistently outperforms machine-composed music, indicating the gaps between human-composed and machine-composed music. Loop coherence scores are positive but close to 0 for all generated music groups, indicating loop coherence to listeners is not very strong, and we observe a slight difference in the median loop coherence scores between human and machine groups. The Happy Emotion and Sad Emotion Scores are obtained from results to the emotion question (Question 5) for pieces classified as happy and sad, respectively. The human group achieves the best result in HES and HES-SES difference. The difference obtained for the Happy and Sad pieces for the machine-composed groups (all_information only) indicates that the generated music is successful in varying the emotional expression from sad to happy. The boxplot highlights that participants had difficulty differentiating happy and sad music in the Without Tonal Tension group, but were able to do so in the All Information group. Therefore, Tonal tension likely helped in generating human-perceivable happy and sad music. This is supported by the Wilcoxon Signed Rank Test results, which will be covered later.
Table 7: Friedman test result for the three groups of music in the listening test.
Question $\chi^{2}$(2) p-value
Preference 51.66 6.06e-12
Loop 36.00 1.53e-8
Emotion 22.41 1.36e-5
@@
In order to gain a better understanding of the outcomes of the Likert-scale questions, we conducted a Friedman test among the three groups of music. The results, as presented in Table 7, indicate that the source of generation (human, machine with all information, and machine without tension) has a significant impact on the listener’s preference, loop, and emotional perspective (p<1e-4 for all three questions). We also carried out multiple pairwise comparisons using the Wilcoxon Sign Rank Test with a Bonferroni-corrected $\alpha$ level ($\alpha$/3 =.0167). We found significant differences between the Human and All Information groups for preference (Z=3002.50, p<.01), loop coherence (Z=4179.50, p<.01), and emotion (Z=4236.50, p<.01), confirming that human-produced music outperforms the machine-generated one. Additionally, we were interested in exploring the effect of bar-level features (i.e., Tonal Tension) in the generation process. We did not find significant differences between the All Information (including tonal tension) and Without Tonal Tension groups for the preference (Z=4551.50, p=.42) and loop coherence (Z=6372.50, p=.70) indicating that conditioning based on tonal tension may not contribute to improving preference and loop coherence. However, we found a significant difference between the All Information and Without Tonal Tension groups (Z=3860.50, p<.0167) for emotion showing that adding tonal tension in the conditioning improves the generation of emotion-specific music.